OpenStudy (anonymous):

what is the equation for a circle that has a center at 0,0 and that passes through 5,6?

OpenStudy (anonymous):

Your radius is √61 I assume they're asking for pi*radious^2 since the radius is an irrational number that equals 61 when squared, so \[\pi \times \sqrt{61}^{2}\]

OpenStudy (anonymous):

I used the pathogen theorem to find the radius \[a^{2} \times b^{2} = c^{2}\]